# Harmonized Estimation of Subgroup-Specific Treatment Effects in Randomized Trials: The Use of External Control Data

### v2

@@ 2.2 Bias and variance of the harmonized estimator
We derive the sampling distribution of the harmonized estimator in this setting. We first restrict attention to the case where randomization is balanced, i.e. $n_{\cdot,1}^{(r)}=n_{\cdot,0}^{(r)}$, and $\pi=(1/K,\dots,1/K)$. We will later relax the assumptions of this stylized example. Let $q=\frac{n^{(e)}_{\cdot,0}}{n^{(r)}_{\cdot,0}+n^{(e)}_{\cdot,0}}$ be the proportion of controls in the RCT.
@@
for all subgroups $k=1,\dots,K$ and some $\gamma\in\mathbb{R}$. In the remainder of the paper the definition of the SDM assumption will be slightly generalized (see (23) and (33)). In previous work, we identified systematic distortion mechanisms in collections of datasets in neuro-oncology [Rahman et al. 2023, Ventz et al. 2022].
@@
Figure 1A shows how the bias of $\hat{\theta}^{h}_{1:K}$ converges to 00 as $\lambda\to\infty$ in Scenario 2. The estimator $\hat{\theta}^{h}_{1:K}$ is suboptimal in settings where the bias terms $\gamma_{1:K}$ vary substantially across subgroups, such as Scenario 3. Recall that we introduced the harmonized estimator for inference on the treatment effects in small patient subgroups in disease settings where heterogeneous bias mechanisms across subgroups (i.e., relevant variations of $\gamma_{1:k}$ across subgroups) appear implausible. We emphasize that in Figure 1A, Scenario 3 involves a high level of variability of $\gamma_{k}$ across subgroups to illustrate that when the SDM assumption is markedly violated harmonization produces biased estimates. Figure 1B shows that when the $\gamma_{k}$’s vary moderately across subgroups the estimator $\hat{\theta}^{h}_{1:K}$ has lower root mean squared error (RMSE) than $\hat{\theta}_{1:K}^{(r)}$. The estimate $\hat{\theta}^{h}_{1:K}$ has lower variance than $\hat{\theta}^{(r)}_{1:K}$. Also, the variance of $\hat{\theta}^{h}_{1:K}$ is similar to the variance of $\hat{\theta}^{(r+e)}_{1:K}$.
@@ 2.4 Interval estimation
where $\hat{\theta}^{(r)}_{k}=\bar{Y}_{k,1}^{(r)}-\bar{Y}_{k,0}^{(r)}$. We include this interval for comparison. We simulate data using the same scenarios used for Figure 1. For the harmonized estimator we use $\Sigma=diag\left(\frac{Q_{k,k}}{\pi_{k}}\right)$, and $\left(\hat{\theta}^{(r)},\hat{\theta}^{(r+e)}_{1:K}\right)$ as defined by expressions (3) and (2). To compute the cut distribution we set $\tau^{(A1)}=\text{diag}(10^{4})$ and $\tau^{(A2)}=\text{diag}(10^{4})$. For each scenario we simulated 2,000 trials. We illustrate results for subgroup $k=1$. Figure 3 shows that in Scenarios 1 and 2, in which $\hat{\theta}^{h}_{1}$ is unbiased for $\lambda=\infty$, the intervals have coverage nearly identical to the nominal $1-\alpha$ level (when $\lambda>10$). In contrast in Scenario 3, where the SDM assumption is markedly violated and $\hat{\theta}^{h}_{1}$ is biased, as expected the coverage is far below the nominal value (close to 0 %). The harmonized and cut intervals centered on $\hat{\theta}^{h}_{1}$ are substantially shorter than the intervals that we compute using only RCT data. The poor coverage in Scenario 3 emphasizes that harmonization is not sufficient to remove the bias of $\hat{\theta}^{(r+e)}_{1:K}$ when the SDM assumption (i.e., $\gamma_{1}=\dots=\gamma_{K}$, equation (9)) is violated. Importantly, data analyses can detect violations of the SDM assumption. Figure 3 suggests that the analyst should consider harmonization if (i) the violation of the SDM assumption based on context-specific considerations is implausible, and (ii) the data do not provide evidence of a violation.
@@ 3 Harmonized estimators and regression models
In this section we discuss harmonized estimators that allow the investigator to account for pre-treatment covariates $X$. The distributions of relevant prognostic variables, such as patient age, may be different in the RCT and EC populations. Recall that the datasets are $D^{(r)}=\left(Y_{i}^{(r)},T_{i}^{(r)},W_{i}^{(r)},X_{i}^{(r)}\right)_{i\leq n^{(r)}}$ for the RCT and $D^{(e)}=\left(Y_{i}^{(e)},T_{i}^{(e)},W_{i}^{(e)},X_{i}^{(e)}\right)_{i\leq n^{(e)}}$ for the EC. Here $X$ includes $d$ pre-treatment covariates, such as demographics. Also, the harmonized estimator $\hat{\theta}_{1:K}^{h}$ is a transformation of (i) $\hat{\theta}_{1:K}^{(r+e)}$, an estimator of $\theta_{1:K}$ (the subgroup-specific treatment effects) based on $D^{(r)}$ and $D^{(e)}$, and (ii) $\hat{\theta}^{(r)}$, an estimator of $\theta=\pi^{\top}\theta_{1:K}$ (the average treatment effect in the randomized population) based only on $D^{(r)}$. The examples in this Section include linear regression, logistic regression, and propensity score weighted logistic regression. In each case we show how to obtain a BD harmonized estimator that is approximately unbiased when the SDM assumption holds. These results are unified in one proposition (see Proposition S2) on weighted generalized linear models (GLMs) in Section S3.6 of the Supplementary Material.
@@ 3.1 Linear model
Figure 4 illustrates some simulations. We simulated data from the linear model (22), with $\gamma_{1:K}$ defined as in Scenarios 1-3 in Figure 1. We generated a single covariate (i.e. $d=1$), $X_{i}^{(r)}\overset{iid}{\sim}N(0,1)$ for $i=1,\dots,n^{(r)}$ in the RCT and $X_{i}^{(e)}\overset{iid}{\sim}N(2,1)$ for $i=1,\dots,n^{(e)}$ in the EC. The vector of covariate values was the same across simulation replicates. The regression coefficient is $\beta=0.5$. Beyond the addition of the covariate $X_{i}$ all simulation details (RCT design, EC design, and treatment effects) are the same as in Figure 1. We compare three estimators: (i) the pooled OLS estimator $\hat{\theta}^{(r+e)}_{1:K}$ based on model (21), (ii) the BD harmonized estimator that we described, using $\lambda=\infty$ and $\Sigma=\text{diag}\left(\nicefrac{{|b_{k}|}}{{\pi_{k}}}\right)$ (“BD harmonized”), and (iii) the RCT-only OLS estimator $\hat{\theta}^{(r)}_{1:K}$ based on the subgroup-specific model (21) excluding the EC data. The harmonized estimator has variance comparable to the pooled estimator and it is more accurate than the RCT-only analysis in Scenarios 1 and 2. In Scenario 3, with $\gamma_{1:K}\not\propto 1_{K\times 1}$, as expected the harmonized estimator inherits the bias of the pooled estimator, although moderately reduced. See Supplementary Figure S2 (Section S2.8) for similar results based on a range of hypothetical sizes $n^{(e)}$ of the external group.
@@ 3.2 Logistic regression
Figure 5 illustrates a simulation study. We generate outcome data from model (32), with $\delta_{1:K}=(\delta,\dots,\delta)$ for $\delta\in[-1,1]$. This is a scenario with systematic bias across subgroups (SDM assumption holds); the Supplementary Material includes a scenario with bias terms $\delta_{1:K}$ that vary across subgroups. We consider $K=5$ subgroups. We generate a single covariate $X_{i}^{(r)}\overset{iid}{\sim}N(0,1)$ in the RCT and $X_{i}^{(e)}\overset{iid}{\sim}N(2,1)$ in the EC, with the regression coefficient $\beta=0.2$. The other details of the simulations are similar to the scenarios used for Figure 4. We compare four estimators of $\theta_{1:K}$: (i) the pooled estimator $\hat{\theta}_{1:K}^{(r+e)}$, using MLE based on model (29); (ii) an RCT-only estimator denoted $\hat{\theta}^{(r)}_{1:K}$, using MLEs based on model (29) without the EC data; (iii) the BD harmonized estimator denoted $\hat{\theta}^{h-B}_{1:K}$, with $\lambda=\infty$ and $\Sigma$ selected using equation (37); and (iv) the VD harmonized estimator denoted $\hat{\theta}^{h-V}_{1:K}$, with $\lambda=\infty$ and $\Sigma$ equal to the estimate of the variance of $\hat{\theta}^{(r+e)}_{1:K}$. Panel A shows that the bias of $\hat{\theta}^{(r+e)}_{1}$ is approximately linear in $\delta_{1}$, and based on this fact harmonization can reduce the bias of $\hat{\theta}^{(r+e)}_{1}$ under the SDM assumption (see equation (33)). Panel B shows that the relative performance of the estimators depends on $\delta_{1}$. Panel C shows that harmonization with $\Sigma$ chosen to reduce bias (BD harmonization) or $\Sigma$ chosen as the estimated variance of $\hat{\theta}^{(r+e)}_{1:K}$ (VD harmonization) produce similar estimates in the simulation scenarios that we considered.
@@ 3.3 Propensity score weighting and harmonization
Alternatively, the analyst can choose $\Sigma$ equal to the estimate of the variance of $\hat{\theta}^{(r+e)}_{1:K}$, using popular sandwich estimators [Lunceford and Davidian 2004] or bootstrap procedures.
@@ 5 Discussion
In this work, we introduced harmonized estimators for subgroup analysis that balance potential efficiency gains from the integration of EC data and the risk of bias due to confounding and other distortion mechanisms. The harmonized estimation procedure has two components: (i) an initial estimator $\hat{\theta}_{1:K}^{(r+e)}$ for the subgroup-specific treatment effects that leverages external control data and (ii) an estimator $\hat{\theta}^{(r)}$ for the average treatment effect in the RCT across subgroups that does not involve the use of external information. The proposed harmonization shrinks the initial subgroup-specific estimates $\hat{\theta}_{1:K}^{(r+e)}$ toward a linear subspace of potential subgroup-specific treatment effects with a prevalence-weighted average concordant with the estimate $\hat{\theta}^{(r)}$ obtained using only the RCT data. Any initial estimator $\hat{\theta}^{(r+e)}_{1:K}$ can be used and we provide strategies for selecting the parameters $\lambda$ and $\Sigma$. For example, throughout this paper we discuss parameterizations chosen with the goal of obtaining harmonized estimates $\hat{\theta}^{h}_{1:K}$ that are unbiased or approximately unbiased when the SDM assumption holds.
@@
Harmonization is a procedure applicable to any subgroup analysis that involves RCT data and ECs, including estimators $\hat{\theta}^{(r+e)}_{1:K}$ beyond those discussed in this paper. In Supplementary Section S3.7, we describe additional simulations in which harmonization improves MSE and reduces bias of subgroup-specific estimates $\hat{\theta}^{(r+e)}_{1:K}$ obtained with (i) a Bayesian causal forest model [Hahn et al. 2020], (ii) a hierarchical model with borrowing of information across subgroups [Jones et al. 2011], and (iii) a propensity score adjusted power prior procedure [Lin et al. 2019]. In these simulations we considered the same scenarios as in Sections 3.1 and 3.2, and computed $\hat{\theta}^{(r+e)}_{1:K}$ (the posterior mean of $\theta_{1:K}$), which was then modified following the proposed harmonization approach. The comparisons of the operating characteristics pre- and post-harmonization remained similar to what we illustrated throughout Section 3. We also point interested readers to Supplementary Section S5, where we discuss the bias of $\hat{\theta}^{(r+e)}_{1:K}$ and $\hat{\theta}^{h}_{1:K}$ in the presence of unmeasured predictors.
@@
See the ancillary Supplementary Material file (PDF) for an extended notation dictionary, all derivations, and additional figures.

### v3

@@ 2.2 Bias and variance of the harmonized estimator
We derive the sampling distribution of the harmonized estimator in this setting. We first restrict attention to the case where randomization is balanced, i.e. $n_{\cdot,1}^{(r)}=n_{\cdot,0}^{(r)}$, and $\pi=(1/K,\dots,1/K)$. We will later relax the assumptions of this stylized example. Let $q=\frac{n^{(e)}_{\cdot,0}}{n^{(r)}_{\cdot,0}+n^{(e)}_{\cdot,0}}$ be the proportion of controls who are in the external data.
@@
for all subgroups $k=1,\dots,K$ and some $\gamma\in\mathbb{R}$. In the remainder of the paper the definition of the SDM assumption will be slightly generalized (see (23) and (33)). The SDM assumption will be relevant to interpret the analyses in this paper, and it holds when the difference between the distributions of the outcomes in the EC group and the RCT control arm, conditional on pre-treatment covariates, is identical across subgroups. These differences can be summarized by comparing conditional means (here and Section 3.1), odds ratios (Section 3.2), or other metrics. In previous work, we identified systematic distortion mechanisms in collections of datasets in neuro-oncology [Rahman et al. 2023, Ventz et al. 2022].
There are several potential causes of systematic distortions. For example, the measurement of the outcomes in the external group might include a systematic shift, in all $K$ subgroups, compared to the measurements in the RCT. In particular, when we consider oncology RCTs, different procedures may be used to measure tumor sizes in the trial patients and the external controls. Similarly, the definition of the first day of follow up, or the schedule of radiographic measurements can vary across studies and cause systematic discrepancies in the assessment of response to therapy and tumor progression [Dabush et al. 2021]. We note that in these settings the experimental therapy might produce positive effects in one or a few of the $K$ subpopulations. Systematic distortions might also arise with subtle differences between the administration of the control therapy in the RCT and the external group, such as different levels of adherence to the schedule of the chemoradiation plan.
@@
Figure 1A shows how the bias of $\hat{\theta}^{h}_{1:K}$ converges to 00 as $\lambda\to\infty$ in Scenario 2. The estimator $\hat{\theta}^{h}_{1:K}$ is suboptimal in settings where the bias terms $\gamma_{1:K}$ vary substantially across subgroups, such as Scenario 3. Recall that we introduced the harmonized estimator for inference on the treatment effects in small patient subgroups in disease settings where heterogeneous bias mechanisms across subgroups (i.e., relevant variations of $\gamma_{1:k}$ across subgroups) appear implausible. We emphasize that in Figure 1A, Scenario 3 involves a high level of variability of $\gamma_{k}$ across subgroups to illustrate that when the SDM assumption is markedly violated harmonization produces biased estimates. Figure 1B shows that when the $\gamma_{k}$’s vary moderately across subgroups ($\Delta_{\gamma}<0.25$) the estimator $\hat{\theta}^{h}_{1:K}$ has lower root mean squared error (RMSE) than $\hat{\theta}^{(r)}$, despite the violation of the SDM assumption. With $\Delta_{\gamma}<0.25$ the violation of this assumption is less marked than in Scenario 3, where the $\gamma_{k}$ parameters are highly heterogeneous ($\Delta_{\gamma}=1$ in Panel A). The estimate $\hat{\theta}^{h}_{1:K}$ has lower variance than $\hat{\theta}^{(r)}_{1:K}$. Also, the variance of $\hat{\theta}^{h}_{1:K}$ is similar to the variance of $\hat{\theta}^{(r+e)}_{1:K}$.
@@ 2.4 Interval estimation
where $\hat{\theta}^{(r)}_{k}=\bar{Y}_{k,1}^{(r)}-\bar{Y}_{k,0}^{(r)}$. We include this interval for comparison. We simulate data using the same scenarios used for Figure 1. For the harmonized estimator we use $\Sigma=diag\left(\frac{Q_{k,k}}{\pi_{k}}\right)$, and $\left(\hat{\theta}^{(r)},\hat{\theta}^{(r+e)}_{1:K}\right)$ as defined by expressions (3) and (2). To compute the cut distribution we set $\tau^{(A1)}=\text{diag}(10^{4})$ and $\tau^{(A2)}=\text{diag}(10^{4})$. For each scenario we simulated 2,000 trials. We illustrate results for subgroup $k=1$. Figure 3 shows that in Scenarios 1 and 2, in which $\hat{\theta}^{h}_{1}$ is unbiased for $\lambda=\infty$, the intervals have coverage nearly identical to the nominal $1-\alpha$ level (when $\lambda>10$). In contrast in Scenario 3, where the SDM assumption is markedly violated and $\hat{\theta}^{h}_{1}$ is biased, as expected the coverage is far below the nominal value (close to 0 %). The harmonized and cut intervals centered on $\hat{\theta}^{h}_{1}$ are substantially shorter than the intervals that we compute using only RCT data. The poor coverage in Scenario 3 emphasizes that harmonization is not sufficient to remove the bias of $\hat{\theta}^{(r+e)}_{1:K}$ when the SDM assumption (i.e., $\gamma_{1}=\dots=\gamma_{K}$, equation (9)) is violated. Importantly, data analyses can detect violations of the SDM assumption (see for example Supplementary Section S3.8). Figure 3 confirms that the analyst should consider harmonization if (i) high quality external control data (see Corrigan-Curay et al. [2018] for explicit criteria) are available, (ii) the violation of the SDM assumption based on context-specific considerations is implausible, and (iii) the datasets do not provide evidence of a violation of the SDM assumption.
@@ 3 Harmonized estimators and regression models
In this section we discuss harmonized estimators that allow the investigator to account for pre-treatment covariates $X$. The distributions of relevant prognostic variables, such as patient age, may be different in the RCT and EC populations. Recall that the datasets are $D^{(r)}=\left(Y_{i}^{(r)},T_{i}^{(r)},W_{i}^{(r)},X_{i}^{(r)}\right)_{i\leq n^{(r)}}$ for the RCT and $D^{(e)}=\left(Y_{i}^{(e)},T_{i}^{(e)},W_{i}^{(e)},X_{i}^{(e)}\right)_{i\leq n^{(e)}}$ for the EC. Here $X$ includes $d$ pre-treatment covariates, such as demographics. Also, the harmonized estimator $\hat{\theta}_{1:K}^{h}$ is a transformation of (i) $\hat{\theta}_{1:K}^{(r+e)}$, an estimator of $\theta_{1:K}$ (the subgroup-specific treatment effects) based on $D^{(r)}$ and $D^{(e)}$, and (ii) $\hat{\theta}^{(r)}$, an estimator of $\theta=\pi^{\top}\theta_{1:K}$ (the average treatment effect in the randomized population) based only on $D^{(r)}$. The examples in this Section include linear regression, logistic regression, and propensity score weighted logistic regression. In each case we show how to obtain a BD harmonized estimator that is approximately unbiased when the SDM assumption holds. These results are unified in one proposition (see Proposition S2) on weighted generalized linear models (GLMs) in Section S3.6 of the Supplementary Material. The analytic expressions of $\Sigma$ for BD harmonization in this Section are special cases of Propositions S3 and S5 in the Supplementary Material.
@@ 3.1 Linear model
Figure 4 illustrates some simulations. We simulated data from the linear model (22), with $\gamma_{1:K}$ defined as in Scenarios 1-3 in Figure 1. We generated a single covariate (i.e. $d=1$), $X_{i}^{(r)}\overset{iid}{\sim}N(0,1)$ for $i=1,\dots,n^{(r)}$ in the RCT and $X_{i}^{(e)}\overset{iid}{\sim}N(2,1)$ for $i=1,\dots,n^{(e)}$ in the EC. The vector of covariate values was the same across simulation replicates. The regression coefficient is $\beta=0.5$. Beyond the addition of the covariate $X_{i}$ all simulation details (RCT design, EC design, and treatment effects) are the same as in Figure 1. We compare three estimators: (i) the pooled OLS estimator $\hat{\theta}^{(r+e)}_{1:K}$ based on model (21), (ii) the BD harmonized estimator that we described, using $\lambda=\infty$ and $\Sigma=\text{diag}\left(\nicefrac{{|b_{k}|}}{{\pi_{k}}}\right)$ (“BD harmonized”), and (iii) the RCT-only OLS estimator $\hat{\theta}^{(r)}_{1:K}$ based on the subgroup-specific model (21) excluding the EC data. The harmonized estimator has variance comparable to the pooled estimator and it is more accurate than the RCT-only analysis in Scenarios 1 and 2. In Scenario 3, with $\gamma_{1:K}\not\propto 1_{K\times 1}$, as expected the harmonized estimator inherits the bias of the pooled estimator, although moderately reduced.
See Supplementary Figure S3 (Section S2.8) for similar comparisons of these estimators that focus on the bias, variance, and RMSE when the trial size $n^{(r)}=100$ is fixed (with $n^{(r)}_{.,0}=33$ and $n^{(r)}_{.,1}=67$), the EC sample size $n^{(e)}$ increases from $33$ to $627$ (i.e., $q=n^{(e)}_{.,0}(n^{(e)}_{.,0}+n^{(r)}_{.,0})^{-1}$ increases from 0.5 to 0.95) and the SDM assumption is violated. The bias and RMSEs of both the pooled and harmonized estimators increase with $n^{(e)}$. As expected this trend is less marked for the harmonized estimator compared to the pooled estimator. Importantly, when $n^{(e)}$ is large other definitions of $\hat{\theta}^{(r+e)}_{1:K}$, beyond pooling (e.g., Bayesian estimators based on hierarchical borrowing or power priors), can be used to reduce the outlined risks associated with distortion mechanisms that violate the SDM assumption.
@@ 3.2 Logistic regression
Figure 5 illustrates a simulation study. We generate outcome data from model (32), with $\delta_{1:K}=(\delta,\dots,\delta)$ for $\delta\in[-1,1]$. This is a scenario with systematic bias across subgroups (SDM assumption holds); the Supplementary Material includes a scenario with bias terms $\delta_{1:K}$ that vary across subgroups. We consider $K=5$ subgroups. We generate a single covariate $X_{i}^{(r)}\overset{iid}{\sim}N(0,1)$ in the RCT and $X_{i}^{(e)}\overset{iid}{\sim}N(2,1)$ in the EC, with the regression coefficient $\beta=0.2$. The other details of the simulations are similar to the scenarios used for Figure 4. We compare four estimators of $\theta_{1:K}$: (i) the pooled estimator $\hat{\theta}_{1:K}^{(r+e)}$, using MLE based on model (29); (ii) an RCT-only estimator denoted $\hat{\theta}^{(r)}_{1:K}$, using MLEs based on model (29) without the EC data; (iii) the BD harmonized estimator denoted $\hat{\theta}^{h-B}_{1:K}$, with $\lambda=\infty$ and $\Sigma$ selected using equation (37); and (iv) the VD harmonized estimator denoted $\hat{\theta}^{h-V}_{1:K}$, with $\lambda=\infty$ and $\Sigma$ equal to the estimate of the variance of $\hat{\theta}^{(r+e)}_{1:K}$. Panel A shows that the bias of $\hat{\theta}^{(r+e)}_{1}$ is approximately linear in $\delta_{1}$, and based on this fact harmonization can reduce the bias of $\hat{\theta}^{(r+e)}_{1}$ under the SDM assumption (see equation (33)). Panel B shows that the relative performance of the estimators depends on $\delta_{1}$. Panel C shows that harmonization with $\Sigma$ chosen to reduce bias (BD harmonization) or $\Sigma$ chosen as the estimated variance of $\hat{\theta}^{(r+e)}_{1:K}$ (VD harmonization) produce similar estimates in the simulation scenarios that we considered. Supplementary Figure S5 illustrates scenarios in which the SDM assumption is violated ($\delta_{1:K}=(\delta-0.5,\phantom{i}\delta+0.5,\phantom{i}\delta-0.5,\phantom{i}\delta+0.5,\phantom{i}\delta-0.5)$ with $\delta\in[-1,1]$) and the harmonized estimator presents lower RMSE than the RCT-only estimator despite post-harmonization bias.
@@ 3.3 Propensity score weighting and harmonization
Alternatively, the analyst can choose $\Sigma$ equal to the estimate of the variance of $\hat{\theta}^{(r+e)}_{1:K}$, using popular sandwich estimators [Lunceford and Davidian 2004] or bootstrap procedures. See Supplementary Figure S3 (Section S2.8) for similar comparisons of these estimators that focus on the bias, variance, and RMSE when the trial size $n^{(r)}=100$ is fixed (with $n^{(r)}_{.,0}=33$ and $n^{(r)}_{.,1}=67$), the EC sample size $n^{(e)}$ increases from $33$ to $627$ (i.e., $q=n^{(e)}_{.,0}(n^{(e)}_{.,0}+n^{(r)}_{.,0})^{-1}$ increases from 0.5 to 0.95) and the SDM assumption is violated. The bias and RMSEs of both the pooled and harmonized estimators increase with $n^{(e)}$. As expected this trend is less marked for the harmonized estimator compared to the pooled estimator. Importantly, when $n^{(e)}$ is large other definitions of $\hat{\theta}^{(r+e)}_{1:K}$, beyond pooling (e.g., Bayesian estimators based on hierarchical borrowing or power priors), can be used to reduce the outlined risks associated with distortion mechanisms that violate the SDM assumption.
@@ 5 Discussion
In this work, we introduced harmonized estimators for subgroup analysis that balance potential efficiency gains from the integration of EC data and the risk of bias due to confounding and other distortion mechanisms. The harmonized estimation procedure has two components: (i) an initial estimator $\hat{\theta}_{1:K}^{(r+e)}$ for the subgroup-specific treatment effects that leverages external control data and (ii) an estimator $\hat{\theta}^{(r)}$ for the average treatment effect in the RCT across subgroups that does not involve the use of external information. The proposed harmonization shrinks the initial subgroup-specific estimates $\hat{\theta}_{1:K}^{(r+e)}$ toward a linear subspace of potential subgroup-specific treatment effects with a prevalence-weighted average concordant with the estimate $\hat{\theta}^{(r)}$ obtained using only the RCT data. Any initial estimator $\hat{\theta}^{(r+e)}_{1:K}$ can be used and we provide strategies for selecting the parameters $\lambda$ and $\Sigma$. For example, throughout this paper we discuss parameterizations chosen with the goal of obtaining harmonized estimates $\hat{\theta}^{h}_{1:K}$ that are unbiased or approximately unbiased when the SDM assumption holds. In Section 2.1 we described how under some regularity conditions this bias-directed (BD) harmonized estimator can be viewed as a modified version of the estimator $\hat{\theta}^{(r+e)}_{1:K}$ obtained by estimating and subtracting the bias $\psi_{1:K}=E\left(\hat{\theta}^{(r+e)}_{1:K}\right)-\theta_{1:K}$ under the SDM assumption. Similarly, BD harmonization of estimators $\hat{\theta}_{1:K}^{(r+e)}$ based on GLMs (Section 3.2) subtracts from $\hat{\theta}^{(r+e)}_{1:K}$ an estimate of $\psi_{1:K}$. Compared to other strategies to reduce or remove bias due to SDMs from subgroup analyses (see Supplementary Section S3.9 for an example), harmonization ensures coherence between subgroup analyses and the primary analysis (i.e., $\pi^{\top}\hat{\theta}^{h}_{1:K}=\hat{\theta}^{(r)}$).
@@
Harmonization is a procedure applicable to any subgroup analysis that involves RCT data and ECs, including estimators $\hat{\theta}^{(r+e)}_{1:K}$ beyond those discussed in this paper. In Supplementary Section S3.7, we describe additional simulations in which harmonization improves MSE and reduces bias of subgroup-specific estimates $\hat{\theta}^{(r+e)}_{1:K}$ obtained with (i) a Bayesian causal forest model [Hahn et al. 2020], (ii) a hierarchical model with borrowing of information across subgroups [Jones et al. 2011], and (iii) a propensity score adjusted power prior procedure [Lin et al. 2019]. In these simulations we considered the same scenarios as in Sections 3.1 and 3.2, and computed $\hat{\theta}^{(r+e)}_{1:K}$ (the posterior mean of $\theta_{1:K}$), which was then modified following the proposed harmonization approach. The comparisons of the operating characteristics pre- and post-harmonization remained similar to what we illustrated throughout Section 3. In these cases it is challenging to analytically derive the bias of $\hat{\theta}^{(r+e)}_{1:K}$, but we can compute BD harmonized estimators. This can be achieved by estimating the bias of $\hat{\theta}^{(r+e)}_{1:K}$ with resampling procedures (see Supplementary Section S3.10). We also point interested readers to Supplementary Section S5, where we discuss the bias of $\hat{\theta}^{(r+e)}_{1:K}$ and $\hat{\theta}^{h}_{1:K}$ in the presence of unmeasured predictors. These results show that various estimators can be harmonized, and that harmonization can improve their operating characteristics. In addition, harmonization can reduce the bias of $\hat{\theta}^{(r+e)}_{1:K}$ in scenarios in which the model used for estimating the treatment effects $\theta_{1:K}$ is misspecified.
@@
See the attached Supplementary Material file (PDF) for an extended notation dictionary, all derivations, and additional figures.